OpenStudy (anonymous):

25+144 squared is 13

OpenStudy (solomonzelman):

squared means to the second power. it's written like x^2 or \[x^2\] what it means is: \[x \times x\] you multiply x times itself 2 times. so 25+144 squared would mean\[(25+144)^2\],\[169^2\cancel{ = }13\] or you meant square root of 144+25, which is \[\sqrt{25+144}\] which equals\[\sqrt{169}\]\[~~->~~~\sqrt{13^2}~~->~~~13\]